A block weighing 100 N and having dimensions 200 mm on an edge is


allowed to slide down an incline on a film of oil having a thickness of 0.05 mm.



If a linear velocity profile in the oil is assumed, what is the terminal


speed of the block?


The viscosity of the oil is 7*10^-2 Ns/m^2


1
Expert's answer
2022-01-23T16:16:49-0500

Explanation & Calculation


  • During the terminal velocity "\small mg\sin \theta=F". The viscous drag is equal to the component of the weight of the block.
  • Moving on, you can use the Newton's law of viscosity for that viscous force to calculate the terminal velocity.

"\qquad\qquad
\begin{aligned}
\small \frac{F}{A}&=\small\mu.\frac{\Delta v}{\Delta d}\\
\small\frac{mg\sin\theta}{(0.2\,m)^2}&=\small0.07\,Ns^2m^{-2}\times\frac{(V_t-0)}{0.00005\,m}
\end{aligned}"

  • You know the weight (mg) of the block and the inclination ("\small\theta") and you can give it a try to find the terminal velocity.